OpenStudy (anonymous):

What is the 35th term of the arithmetic sequence where a1 = 13 and a17 = -83?

OpenStudy (anonymous):

-83 - 13 = -96 -96/16 = -6 so it decreases by -6 every term 13 - (6*34) = -191 -191 is the 35th term

OpenStudy (anonymous):

Thank you, thank you! :)

OpenStudy (anonymous):

you're welcome!